Two blocks of weights W1 = W2 = 10KN rest on a rough inclined plane and are connected by a short piece of string as shown. If the coefficient of friction is μ1 = 0.20 and μ2 = 0.30 respectively: Determine a) The inclination of the plane for which sliding will impend.


1
Expert's answer
2021-04-15T10:32:31-0400

Draw the figure:


Sliding will start when the inclination exceeds angle "\theta".

Forces acting on each block (the lower one is 1 and the upper one is 2):


"1.\space x:f_1+T-W\sin\theta=0,\\
y: W\cos\theta-N=0.\\\space\\
2.\space x:f_2-T-W\sin\theta=0,\\
y: W\cos\theta-N=0."

We also know that "f=\mu N" and see that "W\cos\theta=N", so:


"1.\space \mu_1W\cos\theta+T-W\sin\theta=0,\\
2.\space \mu_2W\cos\theta-T-W\sin\theta=0."

To exclude T, add these equations:


"W(\mu_1+\mu_2)\cos\theta-2W\sin\theta=0,\\
(\mu_1+\mu_2)\cos\theta=2\sin\theta,\\\space\\
\tan\theta=\frac{\mu_1+\mu_2}{2},\\\space\\
\theta=\arctan\frac{\mu_1+\mu_2}{2}=14°."